# Sparse Hyperparametric Itakura-Saito NMF via Bi-Level Optimization

### v2

@@ 5.3 Results on real-world dataset
As shown in Figure 5, the proposed algorithm outperforms other methods in terms of the convergence of the Response function. From Figure 6, we observe that SHINBO achieves the highest ENVSI on $\bm{H}$. A higher value of the ENVSI score (above 0.77) and a lower number of outliers confirm that the proposed algorithm finds the SOI in the observed diagnostic signal, which is consistent with our prior knowledge about the physical state of the tested rolling bearing.
Figure 5: Behavior of the Response function (outer problem) for the real dataset.
Figure 6: The ENVSI score of $\bm{H}$ of the four algorithms: from the left to right are MU with $\lambda=0$, MU with $\lambda=0.1$, MU with $\lambda=0.5$ and SHINBO.
These results are also confirmed by the statistical comparisons. The Kruskall-Wallis tests present $p$-values lesser than the fixed significant level (0.04) and the details of the pair-wise comparison with a BH comparison are presented in Table 4.
@@
The concept is that SHINBO is designed to effectively identify which components of $\bm{\lambda}$ are associated with noise and which are linked to the true sparse signal. The goal is to apply greater penalization to certain components, thereby filtering out the noise while preserving the component representing the SOI. The results indicate that SHINBO performs very well in this context, as it successfully isolates and preserves the meaningful signal while suppressing irrelevant noise.
@@ 6 Conclusion
In this work, we addressed the critical challenge of selecting penalty hyperparameters in NMF by introducing SHINBO, a novel algorithm that employs a bi-level optimization framework to adaptively tune row-dependent penalties. By focusing on the IS divergence, SHINBO proves highly effective for extracting low spectral density components in spectrograms, particularly in the presence of noise. The ability of the algorithm to enforce sparsity constraints and dynamically adjust penalties ensures a more precise separation of meaningful signals from noisy disturbances.
@@
The results highlight SHINBO’s potential to significantly enhance signal recovery in complex, noise-dominated environments. By tackling the hyperparameter selection problem with an adaptive, data-driven approach, SHINBO not only advances the field of NMF but also provides a robust tool for applications requiring precise spectral decomposition and noise suppression. Future work will explore the scalability of SHINBO for larger datasets and its adaptability to other domains with similar challenges.
@@ Acknowledgment
The authors would like to express their sincere appreciation to Professor Radoslaw Zimroz from Faculty of Geoengineering, Mining and Geology at Wroclaw University of Science and Technology, for providing the real data collected in his laboratory.
@@
N.D.B., F.E., and L.S. are partially supported by ”INdAM - GNCS Project”, CUP: E53C24001950001.
N.D.B. and F.E. are supported by Piano Nazionale di Ripresa e Resilienza (PNRR), Missione 4 “Istruzione e Ricerca”-Componente C2 Investimento 1.1, “Fondo per il Programma Nazionale di Ricerca e Progetti di Rilevante Interesse Nazionale”, Progetto PRIN-2022 PNRR, P2022BLN38, Computational approaches for the integration of multi-omics data. CUP: H53D23008870001.

### v3

@@ 5.3 Results on real-world dataset
We evaluate the proposed IS-NMF model with row-wise penalty adaptation with SHINBO on a vibration signal from a faulty rolling bearing (Fig. 2). Following the experimental setup, NMF is applied to the signal spectrogram, and the temporal activations (rows of $\bm{H}$) are shown in Fig. 6(a). One activation (here, the second component) exhibits a sparse, periodic, spike-like pattern characteristic of the SOI, i.e., fault-induced impulsivity in the bearing. Because of NMF’s permutation ambiguity, the SOI can appear in any row; the remaining rows capture latent components associated with broadband and transient disturbances.
Fig. 6(b) displays the envelope spectra of the activations from Fig. 6(a). The first harmonic at approximately 9191Hz aligns with the shaft rotational speed, corroborating that the second component corresponds to the SOI. To quantify spectral leakage and identifiability, the ENVSI metric was used. Fig. 7 summarizes the quantitative results. The right panel shows that SHINBO attains the highest ENVSI, with values above $0.77$ and fewer outliers, indicating reliable SOI recovery consistent with our prior knowledge about the physical state of the tested rolling bearing. The left panel reports the convergence of the Response function, where the proposed method outperforms the baselines.
Refer to caption                 (b)
Figure 6: (a) Time-domain components (the rows of $\bm{H}$) estimated with SHINBO from the faulty signal. (b) Envelope spectrum of the estimated time-domain components. The respective ENVSI scores are given in the subtitles.
Figure 7: Left: Behavior of the Response function (outer problem) for the real dataset. Right: The ENVSI score of $\bm{H}$ of the four algorithms: from left to right are MU with $\lambda=0$, MU with $\lambda=0.1$, MU with $\lambda=0.5$, and SHINBO.
These results are also confirmed by the statistical comparisons. The Kruskal-Wallis tests present $p$-values lesser than the fixed significant level (0.04), and the details of the pairwise comparison with a BH comparison are presented in Table 8, which summarizes the results for the ENVSI coefficient. The $p$-values indicate no significant differences among the MU variants, whereas SHINBO shows statistically significant improvements with respect to all MU configurations, further confirming its effectiveness in capturing relevant structural information.
@@
The concept is that SHINBO is designed to effectively identify which components of $\bm{\lambda}$ are associated with noise and which are linked to the true sparse signal. The goal is to apply greater penalization to certain components, thereby filtering out the noise while preserving the component representing the SOI. The results indicate that SHINBO successfully isolates the meaningful signal while suppressing irrelevant noise.
@@ 6 Conclusion
In this work, we addressed the critical challenge of selecting penalty hyperparameters in NMF by introducing SHINBO, a novel algorithm that employs a bi-level optimization framework to adaptively tune row-dependent penalties. The core idea is that SHINBO does not require setting the penalty coefficient value as a fixed parameter in advance; instead, it automatically optimizes the hyperparameter during the process. Compared to standard algorithms, where the penalty coefficient value is user-dependent, SHINBO consistently achieves high SIR, sparsity, and ENVSI on synthetic and real datasets in a fully automatic manner. This demonstrates that the proposed method can solve the problem in a one-shot way, without relying on any prior parameter tuning, while still providing optimal results in terms of: i)identification: SHINBO yields optimal SIR for both reconstructed matrices; ii) Sparsity: it respects the constraints imposed by the penalizations on the matrix $\mathbf{H}$; iii)Spectral analysis: it ensures an effective analysis of the signal spectrum. By focusing on the IS divergence, SHINBO proves highly effective for extracting low spectral density components in spectrograms, particularly in the presence of noise. The ability of the algorithm to enforce sparsity constraints and dynamically adjust penalties ensures a more precise separation of meaningful signals from noisy disturbances.
@@
Overall, the results highlight SHINBO’s potential to significantly improve signal recovery in complex, noise-dominated environments. By tackling the hyperparameter selection problem with an adaptive, data-driven approach, SHINBO not only advances the field of NMF but also provides a robust tool for applications requiring precise spectral decomposition and noise suppression. Future work will explore the scalability of SHINBO for larger datasets and its adaptability to other domains with similar challenges.
@@ Acknowledgment
The authors thank Professor Radoslaw Zimroz from Faculty of Geoengineering, Mining and Geology at Wroclaw University of Science and Technology, for providing the real data collected in his laboratory.
@@
N.D.B., F.E., and L.S. are partially supported by “INdAM - GNCS Project”, CUP: E53C24001950001. N.D.B. and F.E. are supported by Piano Nazionale di Ripresa e Resilienza (PNRR), Missione 4 “Istruzione e Ricerca”-Componente C2 Investimento 1.1, “Fondo per il Programma Nazionale di Ricerca e Progetti di Rilevante Interesse Nazionale”, Progetto PRIN-2022 PNRR, P2022BLN38, Computational approaches for the integration of multi-omics data. CUP: H53D23008870001.
@@
Declaration of generative AI and AI-assisted technologies
The authors used ChatGPT 5.1 (OpenAI) in order to revise the English grammar. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.